40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4808     Entry time: Mon Jun 13 12:34:21 2011
Author: Jamie, Koji 
Type: Update 
Category: LSC 
Subject: Updated LSC model installed 

After a couple of hickups, I was able to compile and install Koji's rework of the LSC model.

The main changes are that the model now use an RF_PD library part, and the channel names were tweaked to be more in line with what we expect.

I found a couple of small bugs in the model that were preventing it from compiling.  Those were fixed and it compiled with no further problems.

There was also some rearrangement of signal inputs to the PD_DOF matrix.  The matrix screen was updated to reflect the proper inputs.  However, this also meant that the burt restore scripts for the IFO configurations were setting the wrong elements in the matrix.  I fixed the settings for X and Y arm locking, and updated the burt snapshots using the burt/c1ifoconfigure/C1save{X,Y}arm scripts.  NOTE: burt settings will need to be updated for the MICH, PRM, DRM, and FULL IFO configurations as well.

During the build/install process, Joe and I also found a bug in the feCodeGen that was causing the filter screens to be created with the wrong names.  Joe sent out a patch that will hopefully be merged soon.  Building the model with Joe's patch fixed the screen names, so the screens are currently named correctly.

ELOG V3.1.3-